OpenStudy (eva12):

2. Given that f(3)=1,g(3)=-9,f_1(3)=-1,g_1(3)=7 Calculate (f*g)_1(3)=? Calculate (f/g)_1(3)=?

OpenStudy (anonymous):

sorry, what's f_1(3)? I don't get the underscore/underline.

OpenStudy (eva12):

f_1 means the invsere

OpenStudy (eva12):

f^1 inverse

OpenStudy (eva12):

Calculate (f*g)_1(3)=1*-9 Calculate (f/g)_1(3)=1/7

OpenStudy (anonymous):

f_1(3)/g_1(3)=-1/7 F_(3)*g_(3)=-1*7=-7

OpenStudy (anonymous):

f(3)=1, g(3)=-9, f'(3)=-1, g'(3)=7 Calculate (f*g)'= -1 x 7 = -1/7 Calculate (f/g)'= 3/-9 x 7 = -27/7 I'm not sure about this sorry, I'm not use to the way that you have learnt it with the _ , it's too confusing for me.